please guide me on how to solve this problem:
the length of a rectangle is 5cm. more than twice the width. the perimeter of the rectangle is 34 cm. find the dimensions of the rectangle.

Answer by oberobic(2304)

Turn each statement into an equation using the fewest possible unknowns.
.
L = length of the rectangle
.
Length is twice the width plus 5.
L = 2W + 5
.
P = perimeter of a rectangle.
P = 2L + 2W = 2(L + W)
P = 34 (given)
.
Substituting what we know...
34 = 2(L + W)
.
Divide both sides by 2...
17 = L + W
.
Substituting L = 2W+5...
17 = 2W + 5 + W
.
Collecting terms and subtracting 5 from both sides
12 = 3W
so
3W = 12
.
Dividing both sides by 3
W = 4
.
L = 2W+5
L = 2(4)+5
L = 8+5
L = 13
.
Checking the perimeter
P = 2(L+W) = 2(13+4) = 2(17) = 34
.
Checks.